# A real-time digital twin of azimuthal thermoacoustic instabilities

### v1

@@ 4.2 Training the network
During training, the ESN is in open-loop configuration (Figure 3a). The inputs to the reservoir are the training dataset $\bm{\mathsf{U}}=\left[{\bm{\mathit{u}}}_{0}\,|\dots|\,{\bm{\mathit{u}}}_{N_{\mathrm{tr}}-1}\right]$, in which each time component ${\bm{\mathit{u}}}_{k}=[{\bm{\mathit{b}}}_{\mathrm{u}}(t_{k});{{\bm{\mathit{i}}}}_{\mathrm{u}}(t_{k})]$, the subscript ‘u’ indicates training data, and the operator [|][~{}|~{}][ | ] indicates horizontal concatenation. Although we have information from the experimental data to train the network, the optimal parameters of the thermoacoustic system are unknown. Thus, we do not know a priori the model bias and measurement bias. Selecting an appropriate training dataset is key to obtaining a robust ESN, which can estimate the model bias and innovations. We create a set of $L$ guesses on the bias and innovations from a single realization of the experimental data, which means that the ESN is not trained with the ‘true’ bias (which is an unknown quantity in real time). The training data is generated from the experimental data as detailed in C. In summary, the procedure is as follows.
@@
    Take measurements for a training time window $t_{\mathrm{train}}$ of acoustic pressure data $\bm{\mathsf{D}}_{\mathrm{u}}$, and estimate $\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}$ by filtering the power spectral density of $\bm{\mathsf{D}}$ (see §2).
@@
    Generate $L$ model estimates $\bm{\mathsf{Q}}_{\mathrm{u},l}$ of the acoustic pressure from (4) with randomly generated parameters, and correlate in time each $\bm{\mathsf{Q}}_{\mathrm{u},l}$ with $\bm{\mathsf{D}}_{\mathrm{u}}$.
@@
    Compute the training model bias as $\bm{\mathsf{B}}_{\mathrm{u},l}=\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}-\bm{\mathsf{Q}}_{\mathrm{u},l}$ and the training innovations as $\bm{\mathsf{I}}_{\mathrm{u},l}=\bm{\mathsf{D}}_{\mathrm{u}}-\bm{\mathsf{Q}}_{\mathrm{u},l}.$
@@
    Apply data augmentation to improve the network’s adaptivity in a real-time assimilation framework.
The total number of training time series in the proposed training method is $2L$. Training the ESN consists of finding the elements in $\bm{\mathsf{W}}_{\mathrm{out}}$, which minimize the distance between the outputs obtained from an open-loop forecast step of the training data (the features) to the training data at the following time step (the labels). This minimization is solved by ridge regression of the linear system (Lukoševičius, 2012)
@@
where $\lambda$ is the Tikhonov regularization parameter; and $\bm{\mathsf{R}}=\left[[{\bm{\mathit{r}}}_{0};1]\,|\dots|\,[{\bm{\mathit{r}}}_{N_{\mathrm{train}}-1};1]\right]$, with ${\bm{\mathit{r}}}_{0}={\bm{\mathit{0}}}$ and ${\bm{\mathit{r}}}_{k}$ are obtained with (4.1) using the innovations of the training set as inputs. The summations over $2L$ can be performed in parallel to minimize computational costs. Finally, we tune the hyperparameters of the echo state network, i.e., the spectral radius $\rho$, the input scaling $\sigma_{\mathrm{in}}$ and the Tikhonov regularization parameter $\lambda$. We use a recycle validation strategy with Bayesian optimization for the hyperparameter selection (see Racca and Magri, 2021).
@@ 5 Performance of the real-time digital twin
The real-time digital twin has four components: (i) data from sensors (§2.1) (ii) the physics-based low order model (§2.2), (iii) the echo state network to infer the model and measurement biases (§4), and (iv) the real-time and bias-regularized data assimilation method (§3.4) to twin sensors’ data with the low-order model to infer the state and model parameters. We investigate the capability of the proposed real-time digital twin to predict autonomously the dynamics of azimuthal thermoacoustic oscillations. We analyse the digital twin’s performance with respect to established methods: the bias-unregularized ensemble Kalman filter (EnKF) (Nóvoa and Magri, 2022), which is suitable for real-time assimilation; and Langevin-based regression (Indlekofer et al., 2022), which is suitable for offline assimilation. The hyperparameters used to train the ESN, and the assimilation parameters are reported in D.
@@ 5.1 Time-accurate prediction
We focus our analysis on the equivalence ratio $\Phi=0.5125$, which is a thermoacoustically unstable system. Figure 4 shows the time evolution of the acoustic pressure at $\theta=60^{\circ}$. The digital twin is initialized during a short time window up to $0.5$ s. After this, data points from the raw experimental data (red dots) are assimilated every $\Delta t_{\mathrm{d}}=5.86\cdot 10^{-4}$ seconds, which corresponds to approximately $0.64$ data points per acoustic period. At a time instant, the measurement (the data point, red dot) is assimilated into the model, which adaptively updates itself through the r-EnKF and bias estimators (§§3.4,4). After that, the state and model parameters have been updated, the low-order model (7) evolves autonomously until the next data point becomes available. This process mimics a stream of data coming from sensors on the fly. We assimilate 600 measurements during the assimilation window of $0.15$ seconds, the model runs autonomously without seeing more data ($t>0.85$ s).
@@
Although the system has a self-sustained oscillation, the bias-unregularized filter converges towards an incorrect solution, that is, a fixed point. On the other hand, the bias-regularized filter (r-EnKF) successfully learns the thermoacoustic model. Second, the r-EnKF filters out aleatoric noise and turbulent fluctuations. These fluctuations are the difference between the amplitudes of the raw data, which are the input to the digital twin, and the post-processed data, which are the hidden state that we wish to uncover and predict (the presumed acoustic state), of up to approximately 200 Pa (close-ups at the end of assimilation in Figure 4). Third, after assimilation, the r-EnKF predicts the thermoacoustic limit cycle beyond the assimilation window. This means that the digital twin has learnt an accurate model of the system, which generalizes beyond the assimilation. The ensemble model observables ${\bm{\mathit{q}}}_{j}$ and bias-corrected ensemble mean $\bar{{\bm{\mathit{y}}}}=\bar{{\bm{\mathit{q}}}}+{\bm{\mathit{b}}}$ are almost identical at convergence in Figure 4b, which means that the magnitude of the model bias is small (approximately 20 Pa in amplitude, see Figure 7). Fourth, the r-EnKF infers the optimal system’s thermoacoustic parameters, which are seven in this model (Figure 5). The ensemble parameters at $t=0$ in Figure 5 are initialized on physical ranges from the physical parameters used in the literature, which are computed by Langevin regression in (Indlekofer et al., 2022). From inspection of the optimal parameters, we can draw physical conclusions: (i) the parameters of the literature are not the optimal parameters of the model; (ii) the linear growth rate, $\nu$, angular frequency, $\omega$, the heat release strength weighted by the symmetry intensity, $c_{2}\beta$, and the phase of the reactive symmetry, $\Theta_{\epsilon}$, do not significantly vary at regime, which means that these parameters are constants and do not depend on the state; and (iii) the nonlinear parameter, $\kappa$, the amplitude of the reactive symmetry, $\epsilon$, and the direction of asymmetry, $\Theta_{\beta}$ have temporal variations that follow the modulation of the pressure signal envelope (Figure 5). Physically, this means that the optimal deterministic system that represents azimuthal instabilities is a time-varying parameter system (Durbin and Koopman, 2012). By inferring time-varying parameters, we derive a deterministic system that does not require stochastic modelling. We further analyse the thermoacoustic parameters for all the equivalence ratios in Figures 9 and 10.
@@
To summarize, the presence of model biases and measurement biases makes the established bias-unregularized ensemble Kalman filter fail to uncover and predict the physical state from noisy data. In contrast, the bias-regularized filter (r-EnKF) infers the state, parameters, model bias, and measurement bias, which enable a time-accurate and physical prediction beyond the assimilation window.
@@ 5.2 Statistics and biases
On the other hand, the true value of the model bias is not known a priori, but it is presumed as ${\bm{\mathit{b}}}={\bm{\mathit{d}}}^{\dagger}-{\bm{\mathit{q}}}^{\dagger}$, where ${\bm{\mathit{q}}}^{\dagger}$ is the true model estimate, which is not known a priori. The digital twin improves the knowledge that we have on the model bias by correcting the presumed model bias.
@@ 5.3 Physical parameters and comparison with the literature
First, the bias-regularized EnKF infers physical parameters with a small generalization error (the RMS errors are small, depicted with large markers). Large values of the bias regularization, i.e., $\gamma\gtrsim 1$, provide the most accurate set of parameters. This physically means that the low-order model is a qualitatively accurate model because the bias has a small norm (which corresponds to a large $\gamma$). The bias-unregularized EnKF is outperformed in all cases. (The case with $\Phi=0.4875$ is a stable configuration with a trivial fixed point, which is why the parameters have a larger RMS.) Second, we analyse the uncertainty of the parameters, which is shown by an error bar. The height of the error bar is the ensemble standard deviation. The linear growth rate, $\nu$, and the angular frequency, $\omega$, are inferred with small uncertainties. Physically, this means that the model predictions are markedly sensitive to small changes in the linear growth rate and angular frequency. The prediction’s sensitivity to the parameters change with the equivalence ratio. Physically, the large sensitivity of the linear growth rate (i.e., the system’s linear stability) to the operating condition is a characteristic feature of thermoacoustic instabilities, in particular in annular combustors, in which eigenvalues are degenerate (e.g., Magri et al., 2023). Third, there exists multiple combination of parameters that provide a similar accuracy. For example, in the case $\Phi=0.5125$, there are different combinations of $\Theta_{\beta},\epsilon,c_{2}\beta$, which yield a small RMS. Fourth, we compare the digital twin with the state-of-the-art methods based on offline Langevin regression of Indlekofer et al. (2022) (horizontal dashed lines in Figures 9, 10). The inferred parameters are physically meaningful because they lie within value ranges that are similar to the parameters of the literature. The digital twin simultaneously infers all the parameters in real time, which overcomes the limitations of Langevin regression, which needs to be performed on one parameter at a time and offline. The digital twin parameters are optimal because they are the minimizers of (28). The case $\Phi=0.5625$ shows the largest deviation between what the digital twin infers and the results of the offline identification method based on Langevin regression. The solution, which lives on a generalized Bloch sphere (Magri et al., 2023), is further analysed with the quaternion ansätz of Ghirardo and Bothien (2018). Figure 11 shows the acoustic state for $\Phi=0.5625$ from the raw data (panel b), the offline Langevin regression identification (panel c), and the proposed digital twin (panel d). The real-time digital twin infers a set of physical parameters, which correctly capture the physical state of the azimuthal acoustic mode. The nature angle has a bimodal distribution with $|\chi|<\pi/4$, which means that the thermoacoustic instability is a mixed mode. The Langevin regression also provides a bimodal distribution in $\chi$, however, it does not represent the changes in the anti-nodal line, i.e., $\vartheta$. (The parameters identified offline in Indlekofer et al. (2022) and in Figures 9 and 10 were tuned through least-square linear fitting versus $\Phi$. When tuned for a specific $\Phi$, the offline parameters lead to state statistics that more closely match the observables.)
@@
Figure 11: Acoustic state for $\Phi=0.5625$. (a) Bloch sphere for quaternion representation of the acoustic pressure $p(\theta,t)=A\cos{(\vartheta-\theta)}\cos{\chi}\cos{(\omega t+\varphi)}+A\sin{(\vartheta-\theta)}\sin{\chi}\sin{(\omega t+\varphi)}$, where $A$ is the slow-varying amplitude, i.e., the envelope of the acoustic pressure; $\varphi$, is the slow temporal phase drift, $\vartheta$ is the position of the anti-nodal line, and $\chi$ is the nature angle (Ghirardo and Bothien, 2018; Magri et al., 2023). (b) State from the raw data. (c) State probability deduced from Langevin regression methods (Indlekofer et al., 2022). (d) State predicted by the proposed real-time digital twin.

### v2

@@ 4.2 Training the network
During training, the ESN is in open-loop configuration (Figure 4a). The inputs to the reservoir are the training dataset $\bm{\mathsf{U}}=\left[{\bm{\mathit{u}}}_{0}\,|\dots|\,{\bm{\mathit{u}}}_{N_{\mathrm{tr}}-1}\right]$, in which each time component ${\bm{\mathit{u}}}_{k}=[{\bm{\mathit{b}}}_{\mathrm{u}}(t_{k});{{\bm{\mathit{i}}}}_{\mathrm{u}}(t_{k})]$, the subscript ‘u’ indicates training data, and the operator [|][~{}|~{}][ | ] indicates horizontal concatenation. Although we have information from the experimental data to train the network, the optimal parameters of the thermoacoustic system are unknown. Thus, we do not know a priori the model bias and measurement shift. Selecting an appropriate training dataset is key to obtaining a robust ESN, which can estimate the model bias and innovations. We create a set of $L$ guesses on the bias and innovations from a single realization of the experimental data, which means that the ESN is not trained with the ‘true’ bias (which is an unknown quantity in real time). The training data is generated from the experimental data as detailed in C. In summary, the procedure is as follows.
@@
    Take measurements for a training time window $t_{\mathrm{tr}}$ of acoustic pressure data $\bm{\mathsf{D}}_{\mathrm{u}}$, and estimate $\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}$ by applying a Butterworth filter to the raw data $\bm{\mathsf{D}}$ (see §2).
@@
    Generate $L$ model estimates $\bm{\mathsf{Q}}_{\mathrm{u},l}$ of the acoustic pressure from (4). Each $\bm{\mathsf{Q}}_{\mathrm{u},l}$ has a different set of parameters, which are uniformly randomly generated (the parameters’ ranges are reported in D).
@@
    Correlate in time each $\bm{\mathsf{Q}}_{\mathrm{u},l}$ with $\bm{\mathsf{D}}_{\mathrm{u}}$.
@@
    Compute the model bias and innovations as $\bm{\mathsf{B}}_{\mathrm{u},l}=\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}-\bm{\mathsf{Q}}_{\mathrm{u},l}$ and $\bm{\mathsf{I}}_{\mathrm{u},l}=\bm{\mathsf{D}}_{\mathrm{u}}-\bm{\mathsf{Q}}_{\mathrm{u},l}.$
Finally, we apply data augmentation to improve the network’s adaptivity in a real-time assimilation framework. The total number of training time series in the proposed training method is $2L$. Training the ESN consists of finding the elements in $\bm{\mathsf{W}}_{\mathrm{out}}$, which minimize the distance between the outputs obtained from an open-loop forecast step of the training data (the features) to the training data at the following time step (the labels). This minimization is solved by ridge regression of the linear system (Lukoševičius, 2012)
@@
where $\lambda$ is the Tikhonov regularization parameter; and $\bm{\mathsf{R}}=\left[[{\bm{\mathit{r}}}_{0};1]\,|\dots|\,[{\bm{\mathit{r}}}_{N_{\mathrm{tr}}-1};1]\right]$, with ${\bm{\mathit{r}}}_{0}={\bm{\mathit{0}}}$ and ${\bm{\mathit{r}}}_{k}$ are obtained with (4.1) using the innovations of the training set as inputs. The summations over $2L$ can be performed in parallel to minimize computational costs. Finally, we tune the hyperparameters of the echo state network, i.e., the spectral radius $\rho$, the input scaling $\sigma_{\mathrm{in}}$ and the Tikhonov regularization parameter $\lambda$. We use a recycle validation strategy with Bayesian optimization for the hyperparameter selection (see Racca and Magri, 2021).
5 Implementation
To summarize, four stages are necessary to design the real-time digital twin:
1. Initialization:
The ensemble ${\bm{\mathit{\psi}}}_{j}\;\mathrm{for}\;j=1,...,m$, and the ESN are initialized using the parameters reported in Appendix D.
2. Forecast:
Time-march in parallel each ensemble member (21), i.e., the system of thermoacoustic equations and parameters; and the ESN according to (4.1) until observation data become available.
3. Analysis:
Apply the bias-aware r-EnKF (25), which obtains the optimal combination between the unbiased model estimate and the observation data.
4. Re-initialization:
Update the state and parameters of the ensemble with the analyses (i.e., ${\bm{\mathit{\psi}}}^{\mathrm{a}}$ and ${\bm{\mathit{\alpha}}}^{\mathrm{a}}$), and the ESN with the analysis innovation (i.e., $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}={\bm{\mathit{d}}}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}^{\mathrm{a}}$).
Steps (2)-(4) are repeated sequentially as data become available. Once the assimilation process has ended, we forecast further the ensemble and the ESN to analyse the extrapolation and generalization capability on the estimated state, parameters and biases.
@@ 6 Performance of the real-time digital twin
The real-time digital twin has four components: (i) data from sensors (§2.1) (ii) the physics-based low order model (§2.2), (iii) the echo state network to infer the model bias and measurement shift (§4), and (iv) the real-time and bias-regularized data assimilation method (§3.4) to twin sensors’ data with the low-order model to infer the state and model parameters. We investigate the capability of the proposed real-time digital twin to predict autonomously the dynamics of azimuthal thermoacoustic oscillations. We analyse the digital twin’s performance with respect to established methods: the bias-unregularized ensemble Kalman filter (EnKF) (Nóvoa and Magri, 2022), which is suitable for real-time assimilation; and Langevin-based regression (Indlekofer et al., 2022), which is suitable for offline assimilation. The hyperparameters used to train the ESN, and the assimilation parameters are reported in D.
@@ 6.1 Time-accurate prediction
We focus our analysis on the equivalence ratio $\Phi=0.5125$, which is a thermoacoustically unstable system. Figure 5 shows the time evolution of the acoustic pressure at $\theta=60^{\circ}$. Prior to the assimilation, we train the ESN with training data over $t_{\mathrm{tr}}=0.167$ seconds. We start the assimilation at $0.5s$ to avoid using training data during the assimilation, and to have the ensemble in statistically stationary regime. In parallel, we initialize the ESN before the assimilation begins with 10 data points (i.e., the washout). After this, observations from the raw experimental data (red dots) are assimilated every $\Delta t_{d}=5.86\cdot 10^{-4}$ seconds, which corresponds to approximately $0.64$ data points per acoustic period. (The assimilation frequency was selected by down-sampling the raw data, which is recorded at 51.2 kHz, such that $\Delta t_{d}$ fulfils to the Shannon-Nyquist criterion.) At a time instant, the measurement (the data point, red dot) is assimilated into the model, which adaptively updates itself through the r-EnKF and bias estimators (§§3.4,4). After that, the state and model parameters have been updated, the low-order model (7) evolves autonomously until the next data point becomes available. This process mimics a stream of data coming from sensors on the fly. We assimilate 600 measurements during the assimilation window of $0.35s$, the model runs autonomously without seeing more data ($t>0.85$ s).
@@
Although the system has a self-sustained oscillation, the bias-unregularized filter converges towards an incorrect solution, that is, a fixed point. On the other hand, the bias-regularized filter (r-EnKF) successfully learns the thermoacoustic model. Second, the r-EnKF filters out aleatoric noise and turbulent fluctuations. These fluctuations are the difference between the amplitudes of the raw data, which are the input to the digital twin, and the post-processed data, which are the hidden state that we wish to uncover and predict (the presumed acoustic state), of up to approximately 200 Pa (close-ups at the end of assimilation in Figure 5). Third, after assimilation, the r-EnKF predicts the thermoacoustic limit cycle beyond the assimilation window. This means that the digital twin has learnt an accurate model of the system, which generalizes beyond the assimilation. The ensemble model observables ${\bm{\mathit{q}}}_{j}$ and bias-corrected ensemble mean $\bar{{\bm{\mathit{y}}}}=\bar{{\bm{\mathit{q}}}}+{\bm{\mathit{b}}}$ are almost identical at convergence in Figure 5b, which means that the magnitude of the model bias is small (approximately 20 Pa in amplitude, see Figure 8). Fourth, the r-EnKF infers the optimal system’s thermoacoustic parameters, which are seven in this model (Figure 6). The ensemble parameters at $t=0$ in Figure 6 are initialized on physical ranges from the physical parameters used in the literature, which are computed by Langevin regression in (Indlekofer et al., 2022). From inspection of the optimal parameters, we can draw physical conclusions: (i) the parameters of the literature are not the optimal parameters of the model; (ii) the linear growth rate, $\nu$, angular frequency, $\omega$, the heat release strength weighted by the symmetry intensity, $c_{2}\beta$, and the phase of the reactive symmetry, $\Theta_{\epsilon}$, do not significantly vary at regime, which means that these parameters are constants and do not depend on the state; and (iii) the nonlinear parameter, $\kappa$, the amplitude of the reactive symmetry, $\epsilon$, and the direction of asymmetry, $\Theta_{\beta}$ have temporal variations that follow the modulation of the pressure signal envelope (Figure 6). Physically, this means that the optimal deterministic system that represents azimuthal instabilities is a time-varying parameter system (Durbin and Koopman, 2012). By inferring time-varying parameters, we derive a deterministic system that does not require stochastic modelling. We further analyse the thermoacoustic parameters for all the equivalence ratios in Figures 10 and 11.
@@
To summarize, the presence of model bias and measurement shift makes the established bias-unregularized ensemble Kalman filter fail to uncover and predict the physical state from noisy data. In contrast, the bias-regularized filter (r-EnKF) infers the state, parameters, model bias, and measurement shift, which enable a time-accurate and physical prediction beyond the assimilation window.
@@ 6.2 Statistics and biases
On the other hand, the presumed true value of the model bias is not known a priori, but it is presumed as ${\bm{\mathit{b}}}={\bm{\mathit{d}}}^{\dagger}-{\bm{\mathit{q}}}^{\dagger}$, where ${\bm{\mathit{q}}}^{\dagger}$ is the presumed true model estimate, which is not known a priori. The digital twin improves the knowledge that we have on the model bias by correcting the presumed model bias.
@@ 6.3 Physical parameters and comparison with the literature
First, the bias-regularized EnKF infers physical parameters with a small generalization error (the RMS errors are small, depicted with large markers). Large values of the bias regularization, i.e., $\gamma\gtrsim 1$, provide the most accurate set of parameters. This physically means that the low-order model is a qualitatively accurate model because the bias has a small norm (which corresponds to a large $\gamma$). The bias-unregularized EnKF is outperformed in all cases. (The case with $\Phi=0.4875$ is a stable configuration with a trivial fixed point, which is why the parameters have a larger RMS.) Second, we analyse the uncertainty of the parameters, which is shown by an error bar. The height of the error bar is the ensemble standard deviation. The linear growth rate, $\nu$, and the angular frequency, $\omega$, are inferred with small uncertainties. Physically, this means that the model predictions are markedly sensitive to small changes in the linear growth rate and angular frequency. The prediction’s sensitivity to the parameters change with the equivalence ratio. Physically, the large sensitivity of the linear growth rate (i.e., the system’s linear stability) to the operating condition is a characteristic feature of thermoacoustic instabilities, in particular in annular combustors, in which eigenvalues are degenerate (e.g., Magri et al., 2023). Third, there exists multiple combination of parameters that provide a similar accuracy. For example, in the case $\Phi=0.5125$, there are different combinations of $\Theta_{\beta},\epsilon,c_{2}\beta$, which yield a small RMS. Fourth, we compare the digital twin with the offline Langevin regression of Indlekofer et al. (2022) (horizontal dashed lines in Figures 10, 11). The inferred parameters are physically meaningful because they lie within value ranges that are similar to the parameters of the literature. The digital twin simultaneously infers all the parameters in real time, which overcomes the limitations of Langevin regression, which needs to be performed on one parameter at a time and offline. The parameters presented in Figures 10 and 11 depend linearly on $\Phi$. These linear dependencies are approximations of the ground truth and were inferred from the identified set of parameters at each $\Phi$ (see Figure 9 in Indlekofer et al. (2022)). When tuned for a specific $\Phi$, the offline parameters lead to state statistics that more closely match the observables. The digital twin parameters are optimal because they are the minimizers of (24). The long-term statistics of the case $\Phi=0.5625$, which lives on a generalized Bloch sphere (Magri et al., 2023), is further analysed with the quaternion ansätz of Ghirardo and Bothien (2018). Figure 12 shows the acoustic state for $\Phi=0.5625$ from the raw data (panel b), and the proposed digital twin (panel c). The long-term statistics are obtained by forecasting the model (without assimilation) using the identified parameters (reported in E). The real-time digital twin infers a set of physical parameters, which correctly capture the physical state of the azimuthal acoustic mode. The nature angle has a bimodal distribution with $|\chi|<\pi/4$, which means that the thermoacoustic instability is a mixed mode.
@@
Figure 12: Acoustic state for $\Phi=0.5625$. (a) Bloch sphere for quaternion representation of the acoustic pressure $p(\theta,t)=A\cos{(\vartheta-\theta)}\cos{\chi}\cos{(\omega t+\varphi)}+A\sin{(\vartheta-\theta)}\sin{\chi}\sin{(\omega t+\varphi)}$, where $A$ is the slow-varying amplitude, i.e., the envelope of the acoustic pressure; $\varphi$, is the slow temporal phase drift, $\vartheta$ is the position of the anti-nodal line, and $\chi$ is the nature angle (Ghirardo and Bothien, 2018; Magri et al., 2023). (b) State from the raw data. (c) State predicted by the proposed real-time digital twin.
6.4 Generalizability
In this section, we propose a method for generalizing the digital twin to unseen experimental data, e.g., a different equivalence ratio. In sections § 6.1-6.3, we train a different ESN for each equivalence ratio. Here, we test the real-time digital twin framework using an unified ESN, i.e., we train the ESN with data for $\Phi={0.4875,0.5125,0.5375}$ but not for the $\Phi_{\mathrm{test}}=0.5625$ case. The training data generation is identical to that detailed in C, but we combine the training data for $\Phi={0.4875,0.5125,0.5375}$ to train one ESN, which has the same characteristics as those used in the previous analyses (D). We increased the ensemble size to $m=40$ to account for the larger uncertainty and variability in the dynamics. Figure 13 shows the inferred linear growth rate $\nu$ and the heat source strength $c_{2}\beta$ for all $\Phi$. Notably, the digital twin for the equivalence ratio $\Phi_{\mathrm{test}}=0.5625$, which was unseen by the ESN, converges to similar parameters to those in Figure 10. We further analyse the test case $\Phi_{\mathrm{test}}=0.5625$ in Figure 14, which shows the histograms of the acoustic pressure after assimilation. The ESN infers the correct measurement shift to give a zero-mean prediction of the pressure, and a model bias which reduces the distance between the ensemble state estimate and the presumed truth. Overall, we conclude that the unified ESN successfully generalizes to data with unseen dynamics.
Figure 13: Generalizability study. Estimated values of the two stability parameters $c_{2}\beta$ (circles) and $\nu$ (triangles) after data assimilation with the r-EnKF using the same ESN, which is not trained with data for $\Phi=0.5625$. Lighter colours indicate stronger bias regularization parameter; larger marker sizes indicate more accurate solutions, i.e., smaller RMS errors; and the error bars represent the ensemble standard deviation. The dashed lines corresponds to the parameters identified by offline Langevin regression (Indlekofer et al., 2022).
Figure 14: Generalizability study. Histograms of the acoustic pressure at $\Phi=0.5625$ after assimilation with the r-EnKF, at the four observed azimuthal locations. The ESN used in the digital twin is not trained with data for $\Phi=0.5625$. Comparison between the presumed ground truth (black), the observations from raw data (red), the ensemble prediction (filled cyan) and its mean (dashed teal), and bias-corrected ensemble predictions (filled navy) and its mean (dashed light blue). The vertical lines indicate the mean of each of the distributions.